Question 5 Fill in the missing amounts from the folloving T accounts. Accounts Receivable Aug. 10 15 16,80 6,70 Aug. 23 (a) Bal. 8,90 Sept. 5 (b) Sept. 15 8,300 Bal. 4,30 Accounts Payable Aug. 5 (c) 18 3,100 Aug. 29 5,20 Bal. Sept. 12 4,500 7,100 Sept. 23 5,30 Bal. Sales Aug. 10 54,700 Aug. 12 50 15 48,100 Bal. Sept. 12 (f) Sept. 25 40 106,800
Answer #1
The calculation of Aug 23 (a) is shown below
Amount on Aug 10 + Amount on Aug 15 - Balance = Amount of Aug. 23(a)
$16,800 + $6,700 - $8,900 = Amount of Aug. 23(a)
$14,600 Amount of Aug. 23(a)
Since this is the T-account of accounts receivable the transactions of Aug 10 and 15 represents the sales made on credit
and the transaction on Aug 23(a) represents partial payment received from customer which is calculated by adding the sale on credit less the closing balance given
Therefore answer to (a) is $14,600
The calculation of Sept 5 (b) is shown below
Amount on Sept 15 + Ending Balance - Beginning Balance = Amount of Sept. 5 (b)
8,300 + 4,300 - 8,900 = amount of sept 5 (b)
$3,700 Amount of Sept 5 (b)
The transation on Sept 15 represents amount received from customer and on Sept 5 represents the sale on credit made in Sept
which is calculated by adding customer payment with ending balance and deducing opening balance of accounts receivable
Therefore answer to (b) is $3,700
The calculation of Aug 5 (c) is shown below
Amount on Aug 29 + Ending Balance - Amount on Aug 18 = Amount of Aug 5 (c )
5,200 + 4,500 - 3,100 = Amount of Aug 5 (c )
$6,600 Amount of Aug 5 (c )
Since this is the T- account of accounts payable which is created for credit purchases, the transaction on Aug 23 represents payment to the supplier
and on Aug 5 and Aug 18 represents the purchases which are made on credit
Answer to (c ) is $6,600
Ending balance (d)
Beginning Bal + amount on Sept 12 - amount on Sept 23 = End balance
4,500 + 7,100 - 5,300 = End balance
6300 End balance
The transaction on Sept 12 represents credit purchase and that on Sept 23 represents payment made to supplier, the ending balance is calculated as
Beginning bal plus credit purchase less payment to supplier
Answer to (d ) is $6,300
Bal (e )
Amount on Aug 10 + Amount on Aug 15 - Amount on Aug 12 = End balance
54,700 + 48,100 - 500 = End balance
102300 End balance
Since this is the T-account of sales the transactions on 10th and 15th represents the sales made and on 12th represents discount or return, the net sales is calculated as
Sales - return/discount
Answer to (e ) is $102,300
Amount of Sept 12 (f)
Bal + Sept 25 = amount of Sept 12 (f)
106,800 + 400 = amount of sept 12 (f)
107,200 = amount of sept 12 (f)
This is the T-account of sales and the bal represents the net sales of sept which is calculated as sales less the discount/return
Answer to (f ) is $107,200